OpenStudy (anonymous):

I need to write y-64=32(x-4) in standard form

OpenStudy (tkhunny):

Well, do it. Expand those parentheses (Distributive Property) and move things around a little.

OpenStudy (dejavo):

x=1/32y+2

OpenStudy (tkhunny):

Standard Form. Please look this up in your course materials. y-64=32(x-4) y - 64 = 32x - 128 - 64 = 32x - y - 128 64 = 32x - y That's it. OR... 32x - y = 64 You seem to have the arithmetic correct. You just weren't sure where to stop.

OpenStudy (anonymous):

Thank you!